OpenStudy (anonymous):

Solve the system using elimination. 2x + y = 2 2x – y = 6

OpenStudy (anonymous):

Subtract the two equations 2x-2x+y--y=2-6 2y=-4 y=-2 Now go to any equation and subs. y with -2 to get x .

OpenStudy (anonymous):

is there any way you can put that into dummy terms?

OpenStudy (precal):